Question: Mary travels from the city to visit her sister in the country. On the way, she travels 55km/hr. She stays with her sister for 20 hours before starting home. On the return trip, she travels 45 km/hr. She returned home a grand total of 41 hours after she left. How far away, in kilometers, is Mary's house from her sister's?
What I've managed to reason:
41 hours total-20 hours visiting=21 hours traveling
From there, I'm clueless. I've no idea how to start building the equation for this. Help, please?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Mary travels from the city to visit her sister in the country. On the way, she travels 55km/hr. She stays with her sister for 20 hours before starting home. On the return trip, she travels 45 km/hr. She returned home a grand total of 41 hours after she left. How far away, in kilometers, is Mary's house from her sister's?
What I've managed to reason:
41 hours total-20 hours visiting=21 hours traveling
------------
The avg speed of a round trip = 2*55*45/(55+45) km/hr
= 49.5 km/hr
rt distance = r*t = 49.5*21 km
Each way = 49.5*21/2 = 519.75 km
============
The avg speed of a round trip = 2*55*45/(55+45) km/hr
You can trust me on that, or I can prove it.

Answer by stanbon(75887)

Mary travels from the city to visit her sister in the country. On the way, she travels 55km/hr. She stays with her sister for 20 hours before starting home. On the return trip, she travels 45 km/hr. She returned home a grand total of 41 hours after she left. How far away, in kilometers, is Mary's house from her sister's?
-----
City to Country DATA:
rate = 55 kmh ; time = x hrs ; distance = 55x km
------
Country to Cith DATA:
rate = 45 kmh ; time = (21-x) hrs ; distance = 45(21-x) km
---------------------------------------------------------
Equation:
distance = distance
55x = 45(21-x)
55x = 45*21 - 45x
100x = 45*21
x = 9.45 hrs (time going city to country)
---
Distance = 55x = 55*9.45 = 519.75 km
==========================================
Cheers,
Stan H.